OpenStudy (kimberly101):

Use equations to solve the following questions. Imraan is 5 years older than his brother Gareth and the sum of their ages is 31 years. How old is Gareth? (Let x represent Gareth’s age.)

OpenStudy (campbell_st):

let gareths age be x then Imraan is x + 5 then you know the sum of there ages is 31 so x + x+5 = 31 2x + 5 = 31 subtract 5 from both sides 2x = 26 divide by 2 x = 13 gareth is 13 his brother Imraan is 18

OpenStudy (ajprincess):

Let Imraan's age be y. then x+5=y x-y=-5-(1) x+y=31-(2) Add (1) and (2) then u get 2x=26 x=26/2 x=13

OpenStudy (kimberly101):

thankyou so so much i have been trying to do it for so long!!!

OpenStudy (campbell_st):

hope it helps...

OpenStudy (kimberly101):

you helped so much THANKYOU

OpenStudy (kimberly101):

Use equations to solve the following questions David is 5 years younger than his twin brothers. If the sum of their ages is 52, then how old is David?